OpenStudy (anonymous):

convert to a fraction in lowest terms. 20.2% is the answer, 101/500? & 0.22% is the answer, 22/100 =11/50

OpenStudy (anonymous):

Your first answer is correct, the second isn't. Applying the definition of 'percent' to the second, you'll see,\[0.22\times \frac{1}{100}=\frac{22}{100}\times \frac{1}{100}=\frac{22}{10000}=\frac{11}{5000}\]